OpenStudy (anonymous):

Find the area of the triangle with a base of 3 2/3 yd and a height of 3/7 yard

mathslover (mathslover):

Remember that area of a triangle = \(\large{\frac{1}{2}bh}\)

OpenStudy (theviper):

Area of a triangle=1/2*base*height

OpenStudy (theviper):

Put base & height in formula & get ur answer ;)

mathslover (mathslover):

So first of all base = \(\large{3 \frac{2}{3}}\) = \(\large{\frac{(3*2)+2}{3}}\)=\(\large{\frac{8}{3}}\)

OpenStudy (theviper):

Do wht @mathslover says ;)

mathslover (mathslover):

and given height = \(\large{\frac{3}{7}}\) yard put the values .. May i know what r u getting ? @MixMasterTX

OpenStudy (anonymous):

I completing a final exam and this is my last question. However, my reading material converted the height into feet. is this true or is it simply A=1/2(8/3)(3/7)

mathslover (mathslover):

It depends on the options if the options are given in feet then you would have to convert if not then simplify what u got

OpenStudy (theviper):

yes right

OpenStudy (anonymous):

Well they are asking for the answer as ___yd^

OpenStudy (anonymous):

sorry ^2

mathslover (mathslover):

ok so continue to simplify this : A = \(\large{\frac{1}{2}*\frac{8}{3}*\frac{3}{7}}\)

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

just use 1/2 x base x height